I am lost on solving these problems and I need some help
Solve each of the equations. be sure to check the solutions:
square 4x+1 +3 =0
and
square2y+7 + 4=y
thank you

Answer by vertciel(183)

Hello there,
I recategorised your question so you can get a faster answer.
1. 4x + 1 + 3 =0
4x = -4
x = -1
Check: -4 + 1 + 3 = 0

2. square2y+7 + 4=y
I'd like to help you with this but I don't understand what you mean by square2y. Is it the square root of 2y? Please rewrite the question so I can help you.